OpenStudy (anonymous):

What is 1 and 2/3 divided by 2 and 1/6

OpenStudy (praxer):

1 divided by 2 means 1 part out of 2 parts = 1/2 again for 2/3/1/6 we will have = 2/3*6/1=4